# Zo3T: Zero-shot 3D-Aware Trajectory-Guided image-to-video generation via Test-Time Training

### v2

@@ Abstract
Trajectory-Guided image-to-video (I2V) generation aims to synthesize videos that adhere to user-specified motion instructions. Existing methods typically rely on computationally expensive fine-tuning on scarce annotated datasets. Although some zero-shot methods attempt to trajectory control in the latent space, they may yield unrealistic motion by neglecting 3D perspective and creating a misalignment between the manipulated latents and the network’s noise predictions. To address these challenges, we introduce Zo3T, a novel zero-shot test-time-training framework for trajectory-guided generation with three core innovations: First, we incorporate a 3D-Aware Kinematic Projection, leveraging inferring scene depth to derive perspective-correct affine transformations for target regions. Second, we introduce Trajectory-Guided Test-Time LoRA, a mechanism that dynamically injects and optimizes ephemeral LoRA adapters into the denoising network alongside the latent state. Driven by a regional feature consistency loss, this co-adaptation effectively enforces motion constraints while allowing the pre-trained model to locally adapt its internal representations to the manipulated latent, thereby ensuring generative fidelity and on-manifold adherence. Finally, we develop Guidance Field Rectification, which refines the denoising evolutionary path by optimizing the conditional guidance field through a one-step lookahead strategy, ensuring efficient generative progression towards the target trajectory. Zo3T significantly enhances 3D realism and motion accuracy in trajectory-controlled I2V generation, demonstrating superior performance over existing training-based and zero-shot approaches. Code and results are available at https://richard-zhang-ai.github.io/.
@@ 1 Introduction
Recent advances in text- and image-driven video diffusion models have demonstrated remarkable capabilities in generating photorealistic and semantically coherent videos (Ho et al. 2022; Blattmann et al. 2023; Yuan et al. 2025; Chen et al. 2023a; Xu et al. 2025). These developments have paved the way for controllable object motion animation (He et al. 2024a; Namekata et al. 2024; Wang et al. 2025a), which plays a vital role in practical applications such as virtual reality, gaming, advertising, and digital art.
@@
Early works (Wu et al. 2024; Wang et al. 2024b) motion-controllable video generation employed annotated bounding boxes or trajectory points to fine-tune video diffusion models. However, their high computational cost severely limits their practical applicability. In contrast, training-free approaches have attracted significant attention due to their efficiency. These methods modify either the latent representations of target objects through attention mechanisms (Ma, Lewis, and Kleijn 2023; Namekata et al. 2024) or the noise construction (Qiu et al. 2024) during the denoising process, thereby enabling accurate motion control and visually coherent results. However, this paradigm of direct latent space manipulation harbors a fundamental tension. The pre-trained denoising network $\epsilon_{\theta}$ has learned a delicate mapping from a specific distribution of noisy latents to their corresponding noise predictions. Aggressively editing $\mathbf{z}_{t}$ to enforce motion pushes it “off-manifold”, creating a misalignment between the manipulated latent and the model’s learned prior. This model-data misalignment forces the network to denoise a state it is never trained to see, frequently causing a catastrophic degradation in visual quality, manifesting as textural collapse, loss of identity, and other temporal artifacts (He et al. 2024b; Garibi et al. 2024). Moreover, most existing approaches guide motion using 2D trajectories paired with fixed-size bounding boxes or masks. This representation suffers from inherent ambiguity, as it fails to capture the perspective scaling an object should undergo as its depth changes. As shown in Figure 6(a), this lack of 3D awareness leads to physical implausibility and visual distortions, such as unrealistic scaling and motion patterns.
@@
To address these limitations, we propose Zo3T, a novel Test-Time Training (TTT) framework for zero-shot image-to-video (I2V) generation with 3D-aware trajectory control. To mitigate the model-data misalignment in the trajectory guidance process, we introduce a “soft-editing” strategy that co-adapts the data and the model through Trajectory-Guided Test-Time Training. This mechanism injects and optimizes an ephemeral LoRA module within the denoising network at inference time. A trajectory-consistency loss simultaneously updates both the latent state $\mathbf{z}_{t}$ and the temporary LoRA weights. This co-adaptation realigns the model to the edited latent, ensuring the denoising process remains on a stable, high-fidelity manifold while adhering to the specified motion. Building on this, we further introduce Guidance Field Rectification (GFR) via re-evaluating noise scores, which adjusts the guidance field adaptively by maximizing the generation of instance features in the trajectory motion area at each denoising step to adaptively adjust the guidance field. This optimization encourages the latent features to evolve toward the desired trajectory, ensuring controllable and high-quality video synthesis.
To ensure physical plausibility, Zo3T projects user-specified 2D trajectories into a 3D space, leveraging depth cues. This enables perspective-correct motion, ensuring realistic scaling and movement. While recent I2V methods like LeviTor (Wang et al. 2025b) and ObjCtrl-2.5D (Wang et al. 2024a) also incorporate 3D information, LeviTor requires costly training on extensive masked datasets, and ObjCtrl-2.5D’s core assumption of equating object motion with camera movement constrains its applicability in complex scenes. In contrast, our approach offers greater flexibility and directness, requiring no external motion priors or restrictive assumptions.
@@
Figure 2: Advantages of our Method over prior works.
@@ 2 Related Work
The success of diffusion models (Ho, Jain, and Abbeel 2020; Rombach et al. 2022; Song et al. 2021) has shifted the research focus from unconditional generation towards fine-grained, controllable synthesis. While extensively explored in image generation (Zhang, Rao, and Agrawala 2023a; Li et al. 2023; Avrahami et al. 2023), extending such control to the video domain introduces the formidable challenge of maintaining temporal coherence while manipulating spatial content. The dominant approach to this problem has been supervised fine-tuning, where a large, pre-trained video model is adapted to new control modalities on specialized datasets. This paradigm includes methods conditioned on dense signals, such as per-frame pose or depth maps (Zhao et al. 2023; Ma et al. 2023; Chen et al. 2023b), which draw inspiration from image-centric architectures like ControlNet (Zhang, Rao, and Agrawala 2023b).
To offer a more intuitive user interface, a significant lineage of work has focused on sparse trajectory control (Ma, Lewis, and Kleijn 2023; Guo et al. 2023; Zhang et al. 2024). Early attempts often relied on intermediate representations like optical flow (Hao, Huang, and Belongie 2018), which could introduce visual artifacts. More recent methods integrate trajectory guidance directly into the diffusion backbone. For example, DragNUWA (Yin et al. 2023) demonstrated a framework for fusing multi-modal trajectory inputs through a multi-scale architecture. Recognizing the semantic ambiguity of a single moving point, DragAnything (Wu et al. 2024) advanced this concept to the entity-level by leveraging semantic features from the diffusion U-Net to guide the motion of entire objects beyond pixels. Despite their advances, these supervised methods are fundamentally constrained by the high computational cost of retraining and their dependence on large-scale, annotated video datasets, which limits their flexibility and scalability.
@@
Instead of fine-tuning, an alternative approach is to manipulate the generation process at inference, leaving the model weights frozen. This approach was first validated in the image domain (Hertz et al. 2023; Cao et al. 2023; Mou et al. 2024). Inspired by DragGAN (Pan et al. 2023), methods like DragDiffusion (Shi et al. 2024) established that precise, point-based spatial edits could be achieved by directly optimizing the noisy latent code $z_{t}$ with respect to supervisory signals derived from the U-Net’s internal feature maps.
However, translating this zero-shot optimization to video is non-trivial. A key obstacle is the weak temporal correspondence of feature maps within standard video diffusion models (Tang et al. 2023; Luo et al. 2023), which makes them unreliable for guiding motion. An intuitive solution involves a modification to the self-attention mechanism to enforce cross-frame feature alignment (Hertz et al. 2023; Cao et al. 2023; Geyer et al. 2023), thereby creating a usable guidance signal for latent optimization. However, a more fundamental challenge pervades these test-time optimization techniques. The direct, and often aggressive, manipulation of $z_{t}$ can push the latent representation off the manifold learned by the pretrained model (He et al. 2024b; Garibi et al. 2024). This “manifold deviation” disrupts the delicate denoising process, frequently causing a significant degradation in visual quality, manifesting as textural collapse, loss of identity, and other temporal artifacts. Thus, a critical open question remains: how to enforce precise, user-defined control without sacrificing the generative fidelity inherent to the foundational model. Our work directly addresses this tension by introducing a new framework that co-adapts the model and the latent state, ensuring that the guided generation remains on-manifold.
@@ 3 Methodology
This section delineates our framework for zero-shot trajectory-guided video generation. Zo3T is architected to function within the latent space of a pre-trained image-to-video model, such as Stable Video Diffusion (SVD), without requiring any fine-tuning. As illustrated in Figure 3, our method comprises three main components. First, we project user-defined 2D trajectories into a pseudo-3D space, leveraging monocular depth estimation to derive perspective-aware affine transformations. Second, we perform a Test-Time Training (TTT) to co-adapt the latent state and an ephemeral model adapter to enforce feature-space kinematic constraints during the structure-forming stages of denoising. Finally, the Guidance Field Rectification strategy refines the denoising direction via a one-step lookahead optimization, ensuring precise denoising path execution for trajectory control.
@@
Given a well-defined kinematic prior $\mathcal{M}_{k}(\mathbf{p})$ that provides per-frame target regions for guidance, the primary challenge in zero-shot trajectory-guided generation lies in enforcing motion constraints within these regions without disrupting the delicate balance between the latent state $\mathbf{z}_{t}$ and the denoising network $\epsilon_{\theta}$. Prior zero-shot methods (Namekata et al. 2024; Zhang, Alcazar, and Ghanem 2025) often employ “hard edits” to the latents $\mathbf{z}_{t}$ to realize object movement. Such approaches can push the latent state off the learned data manifold into configurations the pre-trained model cannot coherently interpret, resulting in significant generative artifacts.
To resolve this challenge, we introduce a Test-Time Training (TTT) paradigm that performs a “soft adaptation” of both the data and the model. Inspired by previous drag-based control methods (Namekata et al. 2024; Yang et al. 2024; Pan et al. 2023), we leverage the cross-frame similarity of deep features within target regions as trajectory guiding signals. At a specific denoising step $t\in T$, we introduce an ephemeral low-rank perturbation $\Delta\theta_{\text{LoRA}}$ to the model’s parameters and co-adapt it with the latent state $\mathbf{z}_{t}$ by enforcing feature consistency within the target regions. This TTT process seeks a new state-operator equilibrium $(\mathbf{z}_{t}^{*},\theta^{\prime}=\theta\oplus\Delta\theta_{\text{LoRA}})$ while minimizing a test-time objective functional, $\mathcal{J}_{\text{TTT}}$:
@@
Here, $F_{l,k}=\phi_{l}(\mathbf{z}_{t};\theta^{\prime})_{k}$ represents the full feature map from layer $l$ at frame $k$, produced by the adapted model $\theta^{\prime}$. The notation $\mathcal{M}_{b,k}]$ denotes cropping the feature map to to the region defined by the mask for object $b$ at frame $k$. Since the 3D-aware kinematic projection can yields regions of varying sizes across frames, we interpolate all cropped feature maps to a uniform size before computing the pixel-wise loss. The reference features from the first frame, $(F_{l,1})_{\text{frozen}}$, are detached from the computation graph (stop-gradient) to provide a stable optimization target. To mitigate boundary effects, a Gaussian heatmap $\mathcal{G}_{b}$ focuses the loss on the object’s center via the Hadamard product ($\odot$(Namekata et al. 2024). For the selection of feature layers $\mathcal{L}$, we follow the methodology of (Namekata et al. 2024; Wu et al. 2023), which has been proven to yield superior cross-frame spatial-semantic alignment. The TTT process, implemented as a short iterative loop at specific denoising steps, ensures the resulting trajectory is both geometrically accurate and semantically coherent. Trajectory-Guided TTT steers the model, via a semantic loss, to adjust the content layout in subsequent frames, thereby realizing controllable object motion.
@@
While the TTT process successfully constrains the solution space to ensure semantic consistency along the trajectory, it does not prescribe an optimal path within that space. The resulting denoising direction is therefore not guaranteed to be the most efficient or precise for trajectory adherence. Furthermore, the strict spatial consistency enforced by TTT can sometimes lead to unnatural or rigid visual results. To resolve this, we introduce a lookahead-based guidance rectification strategy inspired by the principles of Classifier-Free Guidance (CFG) (Ho and Salimans 2022).
Within the CFG framework, the denoising direction is synthesized from a conditional prediction, which drives the generation towards the control target, and an unconditional prediction, which maintains generative naturality. Their difference can be interpreted as the guidance field, $\mathbf{d}_{t}=\bm{\epsilon}_{t}^{\text{cond}}-\bm{\epsilon}_{t}^{\text{uncond}}$, encapsulating the directional force for control (Li et al. 2025). However, this standard guidance field provides a semantically plausible path of least resistance, which may not align with the globally optimal path for precise trajectory control.
@@
Fidelity Preservation via Fourier Orthogonal Recomposition. While our iterative optimizations achieve effective trajectory control, they risk causing a distributional shift from the diffusion model’s learned prior, leading perceptual quality degradation. Recognizing that motion is predominantly encoded in low-frequency components (Wu et al. 2025), we introduce the Fourier Orthogonal Recomposition (FOR) strategy. This method preserves the high-frequency textural details from the original signal $\mathbf{x}$, which are crucial for maintaining distributional fidelity, while integrating the low-frequency structural modifications from our optimization $\mathbf{x}^{*}$. We define a generalized fusion operator $\mathcal{M}_{fuse}$ as:
@@
where $\mathcal{F}$ is the Fourier Transform and $\mathbf{H}_{\gamma}$ is an ideal low-pass filter mask. This operator is applied twice at the specific denoising step $t$: first to the latent states ($\tilde{\mathbf{z}}_{t}=\mathcal{M}_{fuse}(\mathbf{z}_{t}^{*},\mathbf{z}_{t})$) following TTT, and second to conditional noise predictions ($\tilde{\bm{\epsilon}}_{t,\text{cond}}=\mathcal{M}_{fuse}(\bm{\epsilon}_{t,\text{cond}}^{*},\bm{\epsilon}_{t,\text{cond}})$) after guidance rectification. By consistently sourcing low-frequency structure from our optimizations and high-frequency details from the original signals, this module ensures both control accuracy and high visual fidelity throughout the pipeline.
Selective Timestep Optimization. The coarse structure and motion dynamics of the generated video are largely determined during the early stages of the denoising process (Wang and Vastola 2023). To maximize computational efficiency and ensure structural stability, we apply our trajectory guidance optimizations selectively. Specifically, both the TTT adaptation and GFR are performed only during the early-to-mid denoising phase, from timestep $t=45$ down to $t=30$ in a 50-step schedule. This selective application aligns with established practices in the controllable generation literature (Namekata et al. 2024), striking an effective balance between precise control and generative quality.
@@
DragNUWA (Yin et al. 2023) $126.31$ $251.04$ 10.84 $0.9177$ $0.9272$ $0.5412$ $0.5433$ 320 $\times$ 576 SVD
DragAnything (Wu et al. 2024) $119.07$ $266.42$ $11.64$ $0.9204$ $0.9262$ $0.5469$ $0.5711$ 320 $\times$ 576 SVD
LeviTor (Wang et al. 2025b) $79.41$ $207.23$ $12.06$ $0.9482$ $0.9511$ $0.6417$ $0.6314$ 288 $\times$ 512 SVD
FreeTraj (Qiu et al. 2024) $\checkmark$ $92.12$ $230.21$ $31.63$ $0.9236$ $0.9281$ $0.5841$ $0.5903$ 320 $\times$ 512 VideoCrafter2
ObjCtrl-2.5D (Wang et al. 2024a) $\checkmark$ $81.06$ $212.17$ $18.72$ $0.9512$ $0.9544$ $0.6344$ $0.6429$ 320 $\times$ 576 SVD
SG-I2V (Namekata et al. 2024) $\checkmark$ $79.36$ $209.53$ $14.43$ $0.9448$ $0.9517$ $0.6370$ $0.6317$ 576 $\times$ 1024 SVD

### v3

@@ Abstract
Trajectory-Guided image-to-video (I2V) generation aims to synthesize videos that adhere to user-specified motion instructions. Existing methods typically rely on computationally expensive fine-tuning on scarce annotated datasets. Although some zero-shot methods attempt to trajectory control in the latent space, they may yield unrealistic motion by neglecting 3D perspective and creating a misalignment between the manipulated latents and the network’s noise predictions. To address these challenges, we introduce Zo3T, a novel zero-shot test-time-training framework for trajectory-guided generation with three core innovations: First, we incorporate a 3D-Aware Kinematic Projection, leveraging inferring scene depth to derive perspective-correct affine transformations for target regions. Second, we introduce Trajectory-Guided Test-Time LoRA, a mechanism that dynamically injects and optimizes ephemeral LoRA adapters into the denoising network alongside the latent state. Driven by a regional feature consistency loss, this co-adaptation effectively enforces motion constraints while allowing the pre-trained model to locally adapt its internal representations to the manipulated latent, thereby ensuring generative fidelity and on-manifold adherence. Finally, we develop Guidance Field Rectification, which refines the denoising evolutionary path by optimizing the conditional guidance field through a one-step lookahead strategy, ensuring efficient generative progression towards the target trajectory. Zo3T significantly enhances 3D realism and motion accuracy in trajectory-controlled I2V generation, demonstrating superior performance over existing training-based and zero-shot approaches.
Home Pagehttps://richard-zhang-ai.github.io/
@@ 1 Introduction
Recent advances in text- and image-driven video diffusion models have demonstrated remarkable capabilities in generating photorealistic and semantically coherent videos (ho2022imagen; blattmann2023stable; zhou2025fireedit; yuan2025identity; chen2023videocrafter1; xu2025hunyuanportrait; ma2025followfaster). These developments have paved the way for controllable object motion animation (he2024cameractrl; namekata2024sg; levitor; ma2025followyourmotion), which plays a vital role in practical applications such as virtual reality, gaming, advertising, and digital art (ma2025controllable).
@@
Figure 2: Advantages of our Method over prior works.
@@
Early works (wu2024draganything; wang2024motionctrl; kong2024hunyuanvideo) motion-controllable video generation employed annotated bounding boxes or trajectory points to fine-tune video diffusion models. However, their high computational cost severely limits their practical applicability. In contrast, training-free approaches have attracted significant attention due to their efficiency. These methods modify either the latent representations of target objects through attention mechanisms (ma2023trailblazer; namekata2024sg) or the noise construction (qiu2024freetraj) during the denoising process, thereby enabling accurate motion control and visually coherent results. However, this paradigm of direct latent space manipulation harbors a fundamental tension. The pre-trained denoising network $\epsilon_{\theta}$ has learned a delicate mapping from a specific distribution of noisy latents to their corresponding noise predictions. Aggressively editing $\mathbf{z}_{t}$ to enforce motion pushes it “off-manifold”, creating a misalignment between the manipulated latent and the model’s learned prior. This model-data misalignment forces the network to denoise a state it is never trained to see, frequently causing a catastrophic degradation in visual quality, manifesting as textural collapse, loss of identity, and other temporal artifacts (he2023manifold; ReNoiseInversion). Moreover, most existing approaches guide motion using 2D trajectories paired with fixed-size bounding boxes or masks. This representation suffers from inherent ambiguity, as it fails to capture the perspective scaling an object should undergo as its depth changes. As shown in Figure 6(a), this lack of 3D awareness leads to physical implausibility and visual distortions, such as unrealistic scaling and motion patterns.
@@
To address these limitations, we propose Zo3T, a novel Test-Time Training (TTT) framework for zero-shot image-to-video (I2V) generation with 3D-aware trajectory control. To mitigate the model-data misalignment in the trajectory guidance process, we introduce a “soft-editing” strategy that co-adapts the data and the model through Trajectory-Guided Test-Time Training (Figure 2). This mechanism injects and optimizes an ephemeral LoRA module within the denoising network at inference time. A trajectory-consistency loss simultaneously updates both the latent state $\mathbf{z}_{t}$ and the temporary LoRA weights. This co-adaptation realigns the model to the edited latent, ensuring the denoising process remains on a stable, high-fidelity manifold while adhering to the specified motion. Building on this, we further introduce Guidance Field Rectification (GFR) via re-evaluating noise scores, which adjusts the guidance field adaptively by maximizing the generation of instance features in the trajectory motion area at each denoising step to adaptively adjust the guidance field. This optimization encourages the latent features to evolve toward the desired trajectory, ensuring controllable and high-quality video synthesis. To ensure physical plausibility, Zo3T projects user-specified 2D trajectories into a 3D space, leveraging depth cues. This enables perspective-correct motion, ensuring realistic scaling and movement. While recent I2V methods like LeviTor (levitor) and ObjCtrl-2.5D (objctrl2.5d) also incorporate 3D information, LeviTor requires costly training on extensive masked datasets, and ObjCtrl-2.5D’s core assumption of equating object motion with camera movement constrains its applicability in complex scenes. In contrast, our approach offers greater flexibility and directness, requiring no external motion priors or restrictive assumptions.
@@ 2 Related Work
The success of diffusion models (ho2020denoising; rombach2022high; song2020score) have turned focus from unconditional to fine-grained, controllable synthesis. While extensively explored in image generation (ControlNet; GLIGEN; SpaText; zzz1), extending such control to the video domain introduces the formidable challenge of maintaining temporal coherence while manipulating spatial content (zzz2). The dominant approach to this problem has been supervised fine-tuning, where a large, pre-trained video model is adapted to new control modalities on specialized datasets. This paradigm includes methods conditioned on dense signals, such as per-frame pose or depth maps (zhao2023motiondirector; ma2023follow; chen2023control; jin2024alignment), which draw inspiration from image-centric architectures like ControlNet (ControlNet).
To offer a more intuitive user interface, a significant lineage of work has focused on sparse trajectory control (ma2023trailblazer; guo2023sparsectrl; zhang2024tora; ma2023magicstick). Early attempts often relied on intermediate representations like optical flow (hao2018controllable), which could introduce visual artifacts. More recent methods integrate trajectory guidance directly into the diffusion backbone. For example, DragNUWA (yin2023dragnuwa) demonstrated a framework for fusing multi-modal trajectory inputs through a multi-scale architecture. Recognizing the semantic ambiguity of a single moving point, DragAnything (wu2024draganything) advanced this concept to the entity-level by leveraging semantic features from the diffusion U-Net to guide the motion of entire objects beyond pixels. Despite their advances, these supervised methods are fundamentally constrained by the high computational cost of retraining and their dependence on large-scale, annotated video datasets, which limits their flexibility and scalability.
@@
Instead of fine-tuning, an alternative approach is to manipulate the generation process at inference, leaving the model weights frozen. This approach was first validated in the image domain (PromptToPrompt; MasaCtrl). Inspired by DragGAN (pan2023drag), methods like DragDiffusion (shi2024dragdiffusion) established that precise, point-based spatial edits could be achieved by directly optimizing the noisy latent code $z_{t}$ with respect to supervisory signals derived from the U-Net’s internal feature maps.
However, translating this zero-shot optimization to video is non-trivial. A key obstacle is the weak temporal correspondence of feature maps within standard video diffusion models (tang2023emergent; luo2023dhf), which makes them unreliable for guiding motion. An intuitive solution involves a modification to the self-attention mechanism to enforce cross-frame feature alignment (PromptToPrompt; MasaCtrl; geyer2023tokenflow), thereby creating a usable guidance signal for latent optimization. However, a more fundamental challenge pervades these test-time optimization techniques. The direct, and often aggressive, manipulation of $z_{t}$ can push the latent representation off the manifold learned by the pretrained model (he2023manifold; ReNoiseInversion). This “manifold deviation” disrupts the delicate denoising process, frequently causing a significant degradation in visual quality, manifesting as textural collapse, loss of identity, and other temporal artifacts. Thus, a critical open question remains: how to enforce precise, user-defined control without sacrificing the generative fidelity inherent to the foundational model. Our work directly addresses this tension by introducing a new framework that co-adapts the model and the latent state, ensuring that the guided generation remains on-manifold.
@@ 3 Methodology
Zo3T is architected to function within the latent space of a pre-trained image-to-video model, such as Stable Video Diffusion (SVD), without requiring any fine-tuning. As illustrated in Figure 3, our method comprises three main components. First, we project user-defined 2D trajectories into a pseudo-3D space, leveraging monocular depth estimation to derive perspective-aware affine transformations. Second, we perform a Test-Time Training (TTT) to co-adapt the latent state and an ephemeral model adapter to enforce feature-space kinematic constraints during the structure-forming stages of denoising. Finally, the Guidance Field Rectification strategy refines the denoising direction via a one-step lookahead optimization, ensuring precise denoising path execution for trajectory control.
@@
Given a well-defined kinematic prior $\mathcal{M}_{k}(\mathbf{p})$ that provides per-frame target regions for guidance, the primary challenge in zero-shot trajectory-guided generation lies in enforcing motion constraints within these regions without disrupting the delicate balance between the latent state $\mathbf{z}_{t}$ and the denoising network $\epsilon_{\theta}$. Prior zero-shot methods (namekata2024sg; zhang2025motionawareconceptalignmentconsistent) often employ “hard edits” to the latents $\mathbf{z}_{t}$ to realize object movement. Such approaches can push the latent state off the learned data manifold into configurations the pre-trained model cannot coherently interpret, resulting in significant generative artifacts.
To resolve this challenge, we introduce a Test-Time Training (TTT) paradigm that performs a “soft adaptation” of both the data and the model. Inspired by previous drag-based control methods (namekata2024sg; dav24; pan2023drag), we leverage the cross-frame similarity of deep features within target regions as trajectory guiding signals. At a specific denoising step $t\in T$, we introduce an ephemeral low-rank perturbation $\Delta\theta_{\text{LoRA}}$ to the model’s parameters and co-adapt it with the latent state $\mathbf{z}_{t}$ by enforcing feature consistency within the target regions. This TTT process seeks a new state-operator equilibrium $(\mathbf{z}_{t}^{*},\theta^{\prime}=\theta\oplus\Delta\theta_{\text{LoRA}})$ while minimizing a test-time objective functional, $\mathcal{J}_{\text{TTT}}$:
@@
Here, $F_{l,k}=\phi_{l}(\mathbf{z}_{t};\theta^{\prime})_{k}$ represents the full feature map from layer $l$ at frame $k$, produced by the adapted model $\theta^{\prime}$. The notation $\mathcal{M}_{b,k}]$ denotes cropping the feature map to to the region defined by the mask for object $b$ at frame $k$. Since the 3D-aware kinematic projection can yields regions of varying sizes across frames, we interpolate all cropped feature maps to a uniform size before computing the pixel-wise loss. The reference features from the first frame, $(F_{l,1})_{\text{frozen}}$, are detached from the computation graph (stop-gradient) to provide a stable optimization target. To mitigate boundary effects, a Gaussian heatmap $\mathcal{G}_{b}$ focuses the loss on the object’s center via the Hadamard product ($\odot$(namekata2024sg). For the selection of feature layers $\mathcal{L}$, we follow the methodology of (namekata2024sg; wu2023tune), which has been proven to yield superior cross-frame spatial-semantic alignment. The TTT process, implemented as a short iterative loop at specific denoising steps, ensures the resulting trajectory is both geometrically accurate and semantically coherent. Trajectory-Guided TTT steers the model, via a semantic loss, to adjust the content layout in subsequent frames, thereby realizing controllable object motion.
@@
While the TTT process successfully constrains the solution space to ensure semantic consistency along the trajectory, it does not prescribe an optimal path within that space. The resulting denoising direction is therefore not guaranteed to be the most efficient or precise for trajectory adherence. Furthermore, the strict spatial consistency enforced by TTT can sometimes lead to unnatural or rigid visual results. To resolve this, we introduce a lookahead-based guidance rectification strategy inspired by the principles of Classifier-Free Guidance (CFG) (ho2022cfg).
Within the CFG framework, the denoising direction is synthesized from a conditional prediction, which drives the generation towards the control target, and an unconditional prediction, which maintains generative naturality. Their difference can be interpreted as the guidance field, $\mathbf{d}_{t}=\bm{\epsilon}_{t}^{\text{cond}}-\bm{\epsilon}_{t}^{\text{uncond}}$, encapsulating the directional force for control (li2025flowdirector). However, this standard guidance field provides a semantically plausible path of least resistance, which may not align with the globally optimal path for precise trajectory control.
@@
Fidelity Preservation via Fourier Orthogonal Recomposition. While our iterative optimizations achieve effective trajectory control, they risk causing a distributional shift from the diffusion model’s learned prior, leading perceptual quality degradation. Recognizing that motion is predominantly encoded in low-frequency components (FreeInit), we introduce the Fourier Orthogonal Recomposition (FOR) strategy. This method preserves the high-frequency textural details from the original signal $\mathbf{x}$, which are crucial for maintaining distributional fidelity, while integrating the low-frequency structural modifications from our optimization $\mathbf{x}^{*}$. We define a generalized fusion operator $\mathcal{M}_{fuse}$ as:
@@
where $\mathcal{F}$ is the Fourier Transform and $\mathbf{H}_{\gamma}$ is an ideal low-pass filter mask. This operator is applied twice at the denoising step $t$: first to the latent states ($\tilde{\mathbf{z}}_{t}=\mathcal{M}_{fuse}(\mathbf{z}_{t}^{*},\mathbf{z}_{t})$) following TTT, and second to conditional noise predictions ($\tilde{\bm{\epsilon}}_{t,\text{cond}}=\mathcal{M}_{fuse}(\bm{\epsilon}_{t,\text{cond}}^{*},\bm{\epsilon}_{t,\text{cond}})$) after guidance rectification. By drawing low-frequency structure from optimization and high-frequency details from the original signal, the module preserves control accuracy and visual fidelity.
Selective Timestep Optimization. The coarse structure and motion dynamics of the generated video are largely determined during the early stages of the denoising process (DBLP:journals/corr/abs-2303-02490). To maximize computational efficiency and ensure structural stability, we apply our trajectory guidance optimizations selectively. Specifically, both the TTT adaptation and GFR are performed only during the early-to-mid denoising phase, from timestep $t=45$ down to $t=30$ in a 50-step schedule. This selective application aligns with established practices in the controllable generation literature (namekata2024sg), striking an effective balance between precise control and generative quality.
@@
DragNUWA (yin2023dragnuwa) $126.31$ $251.04$ 10.84 $0.9177$ $0.9272$ $0.5412$ $0.5433$ 320 $\times$ 576 SVD
DragAnything (wu2024draganything) $119.07$ $266.42$ $11.64$ $0.9204$ $0.9262$ $0.5469$ $0.5711$ 320 $\times$ 576 SVD
LeviTor (levitor) $79.41$ $207.23$ $12.06$ $0.9482$ $0.9511$ $0.6417$ $0.6314$ 288 $\times$ 512 SVD
FreeTraj (qiu2024freetraj) $\checkmark$ $92.12$ $230.21$ $31.63$ $0.9236$ $0.9281$ $0.5841$ $0.5903$ 320 $\times$ 512 VideoCrafter2
ObjCtrl-2.5D (objctrl2.5d) $\checkmark$ $81.06$ $212.17$ $18.72$ $0.9512$ $0.9544$ $0.6344$ $0.6429$ 320 $\times$ 576 SVD
SG-I2V (namekata2024sg) $\checkmark$ $79.36$ $209.53$ $14.43$ $0.9448$ $0.9517$ $0.6370$ $0.6317$ 576 $\times$ 1024 SVD